In a quadrilateral OABC, D is the midpoint of BC and E is the point on AB such that AE:ED=2:1 , given that OA=a, OB=b and OC=c express OD and OE in terms of a,b and c


1
Expert's answer
2021-01-29T04:47:42-0500

Given that

"OA:ED=2:1"

"OA\/E=2\/1"

So,

"AB=OB-OA=b-a"

So ,

"AE=AB\/2=(b-a)\/2"

"...Equation 1"

Similarly,

"BC=b-c"

Then,

"CD=(b-c)\/2"

"...... Equation (2)"

"\therefore OD-OC=CD=(b-c)/2"

"\therefore OD-OC=(b-c)/2"

"OD-C=(b-c)\/2"

"OD=C+(b+c)\/2=(b+c)\/2"



"\therefore OD=(b+c)/2"




Again,

"OE-OA=AE"

"OE-OA=(b-a)\/2"

"OE-a=(b-a)\/2"

(Given OA=a)

"OE=(a+b)\/2"






"OE=(a+b)\/2"